OpenStudy (anonymous):

8x + 9y = 9 3x – 5y = 20

OpenStudy (cydney_morgan):

You will need to make one of the pronumerals have an equal positive value and one have an equal negative value so that when you add the two equations together they will cancel each other out: (8x + 9y = 9) * 5 (3x - 5y = 20) * 9 40x + 45y = 45 27x - 45y = 180 Now add the two equations into each other: 67x = 225 x = 225/67 ---------------------------- As x is a fraction it will be very hard to find out y by plugging it in so just do the same as last time: (8x + 9y = 9) * 3 (3x - 5y = 20) * -8 24x + 27y = 27 -24x + 40y = -160 Add the equations into each other: 67y = -133 y = -133/67 -------------------------- x = 225/67 y = -133/67

OpenStudy (anonymous):

(3.36, -2)

OpenStudy (anonymous):

You can give a medal by clicking "Best Response".

OpenStudy (cydney_morgan):

.____________.